# Feature Selection from Differentially Private Correlations

### v1

@@ 4.1. Feature Selection from Private Correlations
Given the performance of SIS in the previous section, our algorithm employs a privatized version of SIS. As a reminder, SIS measures the correlation between each feature and the target variable. The top-$k$ absolutely correlated features are retained. To privately find the top-$k$ highest absolutely correlated features with a target variable, we need to compute private correlations. To do this, we need to bound the sensitivity of the dot product between a feature and the target vector. Unlike the non-private SIS algorithm, which centers and normalizes each feature to a variance of 1, we bound the sensitivity by centering and normalizing each feature to a maximum absolute value of 1. One reason we choose to privatize SIS is because it is easy to understand and can be easily adapted.
@@
1:Design matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$, target vector $\mathbf{y}\in\mathbb{R}^{N}$, privacy parameter $\epsilon$.
2:Center each column in $\mathbf{X}$ and center $\mathbf{y}$. This ensures that each $\mathbf{x}_{(i)}^{\top}\mathbf{y}$ represents a correlation.
3:Scale each column in $\mathbf{X}$ so that the maximum absolute value of any element in each column vector is 1. Scale $\mathbf{y}$ so that the maximum absolute value of any element in the vector is 1. This ensures that the sensitivity of $\lvert\mathbf{x}_{(i)}^{\top}\mathbf{y}\rvert$ is 1.
4:Employ a private top-$k$ selection strategy with parameter $\epsilon$ given that the sensitivity of $\lvert\mathbf{x}_{(i)}^{\top}\mathbf{y}\rvert$ is 1. The “score” or “quality” of feature $i$ is measured with the absolute correlation of the feature with the target, namely $\lvert\mathbf{x}_{(i)}^{\top}\mathbf{y\rvert}$. The private top-$k$ algorithm returns $k$ indices, corresponding to the selected features.
@@ 5. Experiments
The synthetic dataset served as a useful benchmark as it provided a model with a known solution and therefore a known set of top-k features. For the real datasets, we used used Alon (Alon et al., 1999), Borovecki (Borovecki et al., 2005), Burczynski (Burczynski et al., 2006), Chiaretti (Chiaretti et al., 2004), Chin (Chin et al., 2006), Chowdary (Chowdary et al., 2006), Christensen (Christensen et al., 2009), Golub (Golub et al., 1999), Gordon (Gordon et al., 2002), Gravier (Gravier, Eleonore et al., 2010), Khan (Khan et al., 2001), Shipp (Shipp et al., 2002), Singh (Singh et al., 2002), Sorlie (Sørlie et al., 2001), Subramanian (Subramanian et al., 2005), Tian (Tian et al., 2003), West (West et al., 2001), and Yeoh (Yeoh et al., 2002) datasets. These are summarized in Table 1.

### v2

@@ 4.1. Feature Selection from Private Correlations
Given the performance of SIS in the previous section, our algorithm employs a privatized version of SIS. As a reminder, SIS measures the correlation between each feature and the target variable. The top-$k$ absolutely correlated features are retained. To privately find the top-$k$ highest absolutely correlated features with a target variable, we need to compute private correlations. To do this, we need to bound the sensitivity of the dot product between a feature and the target vector. Unlike the non-private SIS algorithm, which centers and normalizes each feature to a variance of 1, we require that the infinity-norm of each column in the dataset is 1 and the infinity-norm of the target vector is 1.111In order to have a better approximation for private correlations, the columns in the dataset must also be centered. This is a common preprocessing technique in regression, and prior works on private regression have assumed datasets with bounded infinity-norms and centered features, see many in (Khanna et al., 2024). We also operate in this setting, but we do recognize that performing this preprocessing in a private manner can be challenging. In our experience, a lot of privacy engineering teams will perform preprocessing in a nonprivate manner, despite this not being ideal. Very recent work has attempted to tackle preprocessing in a private manner, see (Hu et al., 2024). One reason we choose to privatize SIS is because it is easy to understand and can be easily adapted.
@@
1:Design matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ with infinity-norm of each column at most 1, target vector $\mathbf{y}\in\mathbb{R}^{N}$ with infinity-norm at most 1, privacy parameter $\epsilon$.
2:Employ a private top-$k$ selection strategy with parameter $\epsilon$ given that the sensitivity of $\lvert\mathbf{x}_{(i)}^{\top}\mathbf{y}\rvert$ is 1. The “score” or “quality” of feature $i$ is measured with the absolute correlation of the feature with the target, namely $\lvert\mathbf{x}_{(i)}^{\top}\mathbf{y\rvert}$. The private top-$k$ algorithm returns $k$ indices, corresponding to the selected features.
@@ 5. Experiments
The synthetic dataset served as a useful benchmark as it provided a model with a known solution and therefore a known set of top-k features. For the real datasets, we used used Alon (Alon et al., 1999), Borovecki (Borovecki et al., 2005), Burczynski (Burczynski et al., 2006), Chiaretti (Chiaretti et al., 2004), Chin (Chin et al., 2006), Chowdary (Chowdary et al., 2006), Christensen (Christensen et al., 2009), Golub (Golub et al., 1999), Gordon (Gordon et al., 2002), Gravier (Gravier, Eleonore et al., 2010), Khan (Khan et al., 2001), Shipp (Shipp et al., 2002), Singh (Singh et al., 2002), Sorlie (Sørlie et al., 2001), Subramanian (Subramanian et al., 2005), Tian (Tian et al., 2003), West (West et al., 2001), and Yeoh (Yeoh et al., 2002) datasets. These are summarized in Table 1.333Note that we scale and center our synthetic dataset to match the conditions of DP-SIS. Many of the real datasets are already scaled and centered; for those which were not, we performed this operation.